OpenStudy (anonymous):

(8x^3+4x^2+100)/(2x+5)

OpenStudy (ash2326):

We have \[ \frac{(8x^3+4x^2+100)}{(2x+5)}\] We need to find the factors of \(8x^3+4x^2+100\), once we have the factors we'll have an easy road. Are you here?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

We have \[8x^3+4x^2+100\] Generally cubic equations are not straight forward. But I have a feeling that we have one factor as 2x+5, as it's in the denominator. Let's check if I think right. if 2x+5 is a factor then x=-5/2 will make \(8x^3+4x^2+100=0\) Let's check \[8x^3+4x^2+100\] x=-5/2 \[8 ( -5/2)^3+4 (-5/2)^2+100\] \[8*(-125/8)+4(+25/4)+100\] \[-125+25+100=0\] Wow I though right, so we have one factor as (2x+5). Did you get till here?????

OpenStudy (anonymous):

i dont think so

OpenStudy (ash2326):

Which part you don't understand?

OpenStudy (anonymous):

wait.....never mind

OpenStudy (ash2326):

Woah!! tell me, I'm here to help you!!!

OpenStudy (anonymous):

its fine i got it

OpenStudy (ash2326):

Okay so shall we proceed now?

OpenStudy (anonymous):

i need help!!

OpenStudy (anonymous):

on this question, i dont understand it